A table of values for the exponential function y=(1/2)^x has an x-increment of one. If the x-values are changed so that they increase by three rather than one, what is the ratio of the successive y-values? I think it will be 1/8 but I'm not sure.
Answer by bucky(2189)

Well, you can be sure ... you are correct.
.
One way to look at it is to start with the given equation:
.
y = (1/2)^x
.
Then let's just start with x = 0 and increment it by 3 each step.
.
When x = 0 then y = (1/2)^0 = 1 ... (any number to the zeroth power equals 1)
.
When x = 3 then y = (1/2)^3 = 1/8
.
When x = 6 then y = (1/2)^6 = 1/64
.
When x = 9 then y = (1/2)^9 = 1/512
.
Each time x increments by 3 you need to multiply the previous term by 1/8 to get the next
term. So you are correct. Good job and keep up the good work.
.